2.A boy rolls a ball on the table that is 0.85m high.If the boy hits the ball at 4m/s.How far the ball will travel from the base point of the table?


1
Expert's answer
2022-03-31T11:48:26-0400
"L=v_0\sqrt{2H/g}\\
=4*\sqrt{2*0.85/9.8}=1.7\:\rm m"